# Bayesian High-dimensional Linear Regression with Sparse Projection-posterior

### v2

@@ 1 Introduction
High-dimensional linear regression models with numerous predictors, potentially exceeding the observations in number, have garnered significant research interest. Most predictors remain inactive in such models, leading to sparsity in the regression coefficient vector. This feature enables reliable estimation of coefficients by leveraging the underlying low-dimensional structure. Introducing a penalty function to the objective function is a common approach to dealing with such problems, encouraging the minimizer to yield sparser solutions. The popular method LASSO (Tibshirani, 1996; Zhang and Huang, 2008) enforces an $\ell_{1}$-norm constraint on the coefficient vectors, resulting in exact zeros at some coordinates. Variations of the LASSO include the Minimax Concave Penalty (MCP) (Zhang, 2010), Smoothly Clipped Absolute Deviation (SCAD) (Fan and Li, 2001), Dantzig selector (Candes and Tao, 2007), adaptive LASSO (Zou, 2006), non-negative garrotte (Breiman, 1995; Yuan and Lin, 2007) estimators. Earlier research explored LASSO’s asymptotic properties, such as consistency and limiting distribution, in fixed-dimensional settings (Fu and Knight, 2000). A bootstrap LASSO was introduced, but Chatterjee and Lahiri (2010) demonstrated that it lacks consistency when one or more components of the coefficient vector are zero.
@@ 2.1 Motivation
This helps bypass the hassle of MCMC, making computations quite efficient. However, in the end, both approaches lead to a data-dependent probability measure on the derived parameter space to make an inference about the parameter. Another advantage of this method is that the mapping can simultaneously be done with different tuning parameters $\lambda$ per the study’s requirement or purpose. A smaller value of $\lambda$ can be used if the main goal is estimation or prediction, whereas a much larger value of the tuning parameter will be needed for consistent model selection. Optimal frequentist properties guide the choice of these maps.
@@ 3 Main Results
Below, we provide results on the contraction of the posterior at the true value of the parameter and consistency of variable selection. The set of assumptions required for the corresponding results is introduced immediately preceding the results. The assumptions made for the projection-posterior approach are closely related to those used to derive the convergence properties of the LASSO. It should be kept in mind that the assumptions can be changed as per the choice of the penalty for similar posterior convergence results to hold if one chooses to work with any other projection map.
@@ 3.1.1 Estimation and Prediction Consistency
Spike-and-slab or continuous shrinkage posterior may not need all the conditions our method needs for their convergence properties. This is analogous to the conditions needed for an $\ell_{0}$-penalized estimation versus its convex relaxation with the $\ell_{1}$-penalty giving rise to the LASSO. Spike-and-slab or its variants can be considered Bayesian analogs of $\ell_{0}$-penalization methods, requiring exploring all possible models. There is a trade-off between exponential computing complexity versus a few additional restrictions. Moreover, the sparse projection-posterior can be justified by frequentist yardsticks, while most other Bayesian alternatives do not have similar justifications yet.
@@ 4 Numerical Results
We consider the linear regression model with response $\bm{Y}\in\mathbb{R}^{p}$ and fixed design $\bm{X}\in\mathbb{R}^{n\times p}$, and its rows are sampled from $\mathcal{N}_{p}(\bm{0},\bm{\Sigma})$. For the Bayesian method, we use the median probability model (MPM) (Barbieri and Berger, 2004), which chooses those variables in the final model whose marginal posterior probability of being included is at least $0.5$ to select predictors. We use an estimator with zeroes at all components not selected in more than 50% MCMC iterations. The estimator takes the average of all the $R$ samples for the remaining components, where $R$ is the number of MCMC draws. A comparison between Bayesian and frequentist methods included in the study of estimation accuracy is based on the $\mathrm{MSE}=(\hat{\bm{\theta}}-\bm{\theta}^{0})^{\mathrm{T}}\bm{\Sigma}(\hat{\bm{\theta}}-\bm{\theta}^{0}).$ For a variable selection procedure, we define True Positives (TP) as the signals correctly identified to be non-null, False Positives (FP) as the noise variables falsely selected as active variables, True Negatives (TN) as the null variables correctly not selected by the model and finally False Negatives (FN) as the signals mistakenly left out of the model. To compare different variable selection methods, we report three quantities, namely, True Positive Rate (TPR) which is the ratio of correctly selected predictors to the total number of true signals ($s_{0}$), TPR = TP/(TP + FN), False Discovery Proportion (FDP) which is the ratio of falsely selected predictors to the number of selected signals, FDP = FP/(FP+TP) and Matthew’s correlation coefficient (MCC):
@@ 4.1.3 Simulation Results
We investigated the time taken for one run of each method in this study for both small-$p$ and large-$p$ cases. In Table 1 reported are the mean times (in seconds) to finish $R=10000$ MCMC runs of each of the Bayesian methods along with one run of the debiased LASSO and the bootstrap LASSO with $5000$ replicates. The projection method and the bootstrap LASSO have similar computing costs. Computation times under $p=2000$ for SSVS and BLASSO are excessive (in hours) and have not been reported along with the other faster computing methods.
@@
We see in Figure 1 that the MSEs in the low-dimensional model with $n=1000>p=100$ are much smaller than the ones from the high-dimensional model with $p=2000>n=100$ for all the methods and for both design types. Performance-wise, in the low-dimensional setting, the LASSO has the largest estimation error, but other methods have very comparable performance. The Bayesian LASSO was skipped in the high-dimensional setting due to its high computing time. The LASSO and the sparse projection-posterior mean have somewhat higher MSE than other methods for independent predictors, but the difference is minimal for correlated predictors. In all cases, the MCP, horseshoe, SSVS, and SSLASSO generally have similar average estimation errors.
Figure 1: Boxplots of the $n*$MSEs for different methods averaged over $M=100$ replications corresponding to the different cases are provided starting from large-$p$ independent and correlated design, followed by the same for small-$p$.
@@
Projection 1 0.084 (0.063) 0.944 (0.108) 1 0.018 (0.037) 0.991 (0.019)
$n=1000$ $p=100$ $s_{0}=10$ LASSO 1 0.140 (0.013) 0.935 (0.081) 1 0.129 (0.012) 0.979 (0.019)
MCP 1 0.125 (0.013) 0.964 (0.017) 1 0.108 (0.011) 0.996 (0.084)
Horseshoe 1 0.006 (0.023) 0.986 (0.013) 1 0.004 (0.034) 0.991 (0.007)
@@
In Table 2, we report the measures for variable selection corresponding to both the small-$p$ ($n=1000,p=100$) and large-$p$ ($n=100,p=2000$) regimes and both design types. The SSVS maintained its best selection performance in low- and high-dimensional cases, followed by the SSLASSO and the projection method. However, the SSLASSO fails to maintain its consistency under correlated predictors, whereas the projection-posterior method is hardly affected by it. When $p<n$, the Bayesian LASSO selects many noise variables, causing its FDP to be the highest and, consequently, its MCC to be the lowest. The MCP does a good job of identifying the correct signals. The same conclusion can be made from the plots in the third and fourth panels of Figure 2, which show the boxplots of the MCC values of the competing methods. On average, the sparse projection-posterior-based method is comparable to the horseshoe method but performs better when the design matrix has an underlying correlation structure. Overall, from this limited simulation study, it can be concluded that the method proposed here has good selection properties. In addition, as will be shown next, the method can provide coverage for the signals, but the SSVS and SSLASSO fail to do so.
Figure 2: The boxplots of the $n*$MSEs of all the comparing methods are provided for the independent design (first) and the correlated design (second). The boxplots of the MCC values of all the methods are provided for the independent design (third) and the correlated design (fourth). All the plots correspond to $n=p=300$ and $M=100$ replications.
@@
We compare the coverage of 95% confidence or credible regions from the different methods included in the simulation. In Tables 3 and 4, we report the mean coverage averaged over the $M$ replicates corresponding to the uncorrelated and correlated designs respectively, but separately for active and inactive variables. In general, coverage and length-wise, all methods do better when the variables are independent. On the Bayesian side, only the proposed sparse projection-posterior method reaches close to the desired coverage. The average lengths of all the intervals increase significantly to provide the same extent of coverage when the dimension $p$ grows with the sample size $n$. The existing Bayesian methods provide coverage if the parameter is zero but fail to attain 95% coverage for non-zero parameter values corresponding to active predictors. Also, the average coverage of signals deteriorates in the correlated design setting. However, when $p\gg n$, the only Bayesian method that provides good coverage of the signals is the sparse projection-posterior method. The plots in Figure 3 show the coverage and length of credible/confidence regions averaged over all the signal variables for all competing methods. For small-$p$ and $n=p$ cases, the sparse projection posterior method attains desirable coverage with a very small length, making it the best candidate. Its performance degrades when the dimension exceeds $n$ but is still better than most methods.
@@
Projection 0.908 (0.094) 0.453 (0.023) 0.967 (0.004) 0.427 (0.003)
@@
Horseshoe 0.891 (0.099) 0.102 (0.015) 0.999 (0.000) 0.205 (0.005)
Projection 0.945 (0.011) 0.159 (0.006) 0.977 (0.001) 0.267 (0.015)
@@
Projection 0.966 (0.111) 0.103 (0.003) 0.949 (0.030) 0.088 (0.002)
@@
Projection 0.868 (0.131) 0.695 (0.011) 0.982 (0.002) 0.548 (0.007)
@@
Horseshoe 0.845 (0.107) 0.541 (0.029) 0.996 (0.001) 0.361 (0.007)
Projection 0.947 (0.013) 0.315 (0.010) 0.968 (0.007) 0.464 (0.007)
@@
Projection 0.952 (0.101) 0.168 (0.002) 0.969 (0.019) 0.092 (0.008)
@@
In Figure 4, we present individual coverage and interval lengths for two randomly selected signals and two randomly selected noise components in all four settings. These settings compare intervals produced by the horseshoe, sparse projection-posterior, debiased LASSO (ZnZ), and bootstrap LASSO for active (1 and 6) and noise (50 and 100) variables. Both signals have a true coefficient of 2. The plots show frequentist methods, and the Bayesian sparse posterior-projection method offers good coverage and comparable interval lengths. The horseshoe prior, in particular, exhibits slightly shorter lengths, especially for noise variables, consistent with results in Table 3. Importantly, our proposed method consistently covers the true signal.
We also verify the coverage of the credible ellipsoid introduced in Section 3.2.2 for the three cases studied in this section, namely $n=1000,p=100$, $n=p=300$ and finally $n=100,p=1000$. The reported results are the mean coverage values over $M=100$ repetitions. For the low-dimensional case, the average coverage of the credible ellipsoid for the signals is $0.954$ with standard error of $0.004$. When sample size and dimension are both equal to 300, the average coverage and standard error are respectively $0.951$ and $0.008$. Finally, in the higher dimensional setting, we observe the average coverage by the ellipsoid to be $0.947$ with a standard error of $0.010$.
@@
4.2 Simulation Study 2: Distributed Computing and Credible Ellipsoid

### v3

@@ 1 Introduction
High-dimensional linear regression models with numerous predictors, potentially exceeding the observations in number, have garnered significant research interest. Most predictors remain inactive in such models, leading to sparsity in the regression coefficient vector. This feature enables reliable estimation of coefficients by leveraging the underlying low-dimensional structure. Introducing a penalty function to the objective function is a common approach to dealing with such problems, encouraging the minimizer to yield sparser solutions. The popular method LASSO (Tibshirani, 1996; Zhang and Huang, 2008) enforces an $\ell_{1}$-norm constraint on the coefficient vectors, resulting in exact zeros at some coordinates. Variations of the LASSO include the Minimax Concave Penalty (MCP) (Zhang, 2010), Smoothly Clipped Absolute Deviation (SCAD) (Fan and Li, 2001), Dantzig selector (Candes and Tao, 2007), adaptive LASSO (Zou, 2006), non-negative garrotte (Breiman, 1995; Yuan and Lin, 2007) estimators. Earlier research explored LASSO’s asymptotic properties, such as consistency and limiting distribution, in fixed-dimensional settings (Knight and Fu, 2000). A bootstrap LASSO was introduced, but Chatterjee and Lahiri (2010) demonstrated that it lacks consistency when one or more components of the coefficient vector are zero.
@@ 2.1 Motivation
This helps bypass the hassle of MCMC, making computations quite efficient. As a result, we do not have to worry about the convergence of the MCMC and autocorrelation of the obtained samples, and drawing much fewer samples suffices. Owing to the issue of autocorrelated posterior samples in continuous shrinkage methods such as the horseshoe, the Effective Sample Size (ESS) is very low. The ESS for the proposed method, however, will essentially be the same as the number of posterior draws owing to independence. However, in the end, both approaches lead to a data-dependent probability measure on the derived parameter space to make an inference about the parameter. Another advantage of this method is that the mapping can simultaneously be done with different tuning parameters $\lambda$ per the study’s requirement or purpose. A smaller value of $\lambda$ can be used if the main goal is estimation or prediction, whereas a much larger value of the tuning parameter will be needed for consistent model selection. Optimal frequentist properties guide the choice of these maps.
@@
The soft-spike-and-slab prior and continuous shrinkage prior both use immersion maps to induce sparse posteriors on the regression parameter. The first method samples from the posterior using the stochastic search variable selection (SSVS) MCMC, and generates samples from the posterior distribution of $(\theta_{1},\dots,\theta_{p},\gamma_{1},\ldots,\gamma_{p})$, where $\gamma_{1},\ldots,\gamma_{p}$ are variable inclusion indicators. Clearly, the sampled values of the regression parameters are dense and are made sparse by the implicit immersion map $(\theta_{1},\ldots,\theta_{p},\gamma_{1},\ldots,\gamma_{p})\mapsto(\gamma_{1}\theta_{1},\ldots,\gamma_{p}\theta_{p}).$ Thus, in this context, $(\theta_{1},\ldots,\theta_{p},\gamma_{1},\ldots,\gamma_{p})$ are superparameters (as defined in the paper) and $\theta_{j}^{*}=\gamma_{j}\theta_{j}$, $j=1,\ldots,p$, are the parameters of the model. For continuous shrinkage priors, the thresholding map $\theta_{j}^{*}=\theta_{j}\mathbbm{1}\{|\theta_{j}|>\delta_{n}\},j=1,\ldots,p,$ defines the implied immersion map with thresholding parameter $\delta_{n}$, which also needs to be determined and is usually chosen in an ad hoc manner.
The main difference between our method and these is that, by using a slightly more complicated immersion map given by the sparse projection operation, we can avoid MCMC altogether and replace it with independent sampling from a multivariate normal distribution. As a result, we do not have to worry about the convergence of the MCMC and autocorrelation of the obtained samples, and drawing much fewer samples suffices. The sparse projection map is only slightly more challenging to compute, thanks to already developed algorithms using solution paths of the LASSO.
@@ 3 Main Results
Below, we provide results on the contraction of the posterior at the true value of the parameter and consistency of variable selection. The set of assumptions required for the corresponding results is introduced immediately preceding the results. The assumptions made for the projection-posterior approach are closely related to those used to derive the convergence properties of the LASSO. It should be kept in mind that the assumptions can be changed as per the choice of the penalty for similar posterior convergence results to hold if one chooses to work with any other projection map. Moreover, we do not use the Gaussian posterior distribution for any of the proofs. We will only need sub-Gaussianity of $\bm{\eta}$ to address the ultrahigh-dimensional case where $\log p$ is exponential in a power of $n$. For a more modest polynomial growth in $p$, the sub-Gaussianity condition can be significantly weakened to the condition of finiteness of a sufficiently high moment, depending on the relative comparison of $p$ and $n$, provided we use a tuning parameter of a different magnitude.
@@ 3.1.1 Estimation and Prediction Consistency
Spike-and-slab or continuous shrinkage posterior may not need all the conditions our method needs for their convergence properties. This is analogous to the conditions needed for an $\ell_{0}$-penalized estimation versus its convex relaxation with the $\ell_{1}$-penalty giving rise to the LASSO. Spike-and-slab or its variants can be considered Bayesian analogs of $\ell_{0}$-penalization methods, requiring exploring all possible models. There is a trade-off between exponential computing complexity versus a few additional restrictions. The assumptions required for obtaining similar results by the continuous shrinkage prior are more or less similar to our method (Song and Liang, 2017). On the other hand, the modified bootstrap LASSO method, which is consistent and can provide uncertainty quantification for the LASSO estimator assumes fixed $p$. Moreover, the sparse projection-posterior can be justified by frequentist yardsticks, while most other Bayesian alternatives do not have similar justifications yet.
@@ 4 Numerical Results
We consider the linear regression model with response $\bm{Y}\in\mathbb{R}^{p}$ and fixed design $\bm{X}\in\mathbb{R}^{n\times p}$, and its rows are sampled from $\mathcal{N}_{p}(\bm{0},\bm{\Sigma})$. For the Bayesian method, we use the median probability model (MPM) (Barbieri and Berger, 2004), which chooses those variables in the final model whose marginal posterior probability of being included is at least $0.5$ to select predictors. We use an estimator with zeroes at all components not selected in more than 50% MCMC iterations. The estimator takes the average of all the $R$ samples for the remaining components, where $R$ is the number of MCMC draws. A comparison between Bayesian and frequentist methods included in the study of estimation accuracy is based on the $\mathrm{MSE}=(\hat{\bm{\theta}}-\bm{\theta}^{0})^{\mathrm{T}}\bm{\Sigma}(\hat{\bm{\theta}}-\bm{\theta}^{0}).$ For a variable selection procedure, we define True Positives (TP) as the signals correctly identified to be non-null, False Positives (FP) as the noise variables falsely selected as active variables, True Negatives (TN) as the null variables correctly not selected by the model and finally False Negatives (FN) as the signals mistakenly left out by the model. To compare different variable selection methods, we report three quantities, namely, True Positive Rate (TPR) which is the ratio of correctly selected predictors to the total number of true signals ($s_{0}$), TPR = TP/(TP + FN), False Discovery Proportion (FDP) which is the ratio of falsely selected predictors to the number of selected signals, FDP = FP/(FP+TP) and Matthew’s correlation coefficient (MCC):
@@ 4.1.3 Simulation Results
We investigated the time taken for one run of each method in this study for both small-$p$ and large-$p$ cases. In Table 1 reported are the mean times (in seconds) to finish $R=10000$ MCMC runs of each of the Bayesian methods along with one run of the debiased LASSO and the bootstrap LASSO with $5000$ replicates. The projection method and the bootstrap LASSO have similar computing costs. Computation time under $p=2000$ for BLASSO is excessive (in hours) and have not been reported along with the other faster computing methods. Since we reuse the value obtained for the LASSO using cross-validation, and since most of the computational cost is used in tuning, the additional computation cost of our method over the LASSO is not substantial. For instance, running LASSO once with cross-validation for the tuning parameter for a regression problem with $p=2000$ variables, $s_{0}=10$ active variables and $n=100$ sample size is about 0.26 seconds. On the other hand, if we run LASSO repeatedly to project the posterior samples, we take about $\sim 3.6$ seconds for 1000 posterior samples, $\sim 5.2$ seconds for 2000 posterior samples, $\sim 8$ seconds for 3000 posterior samples, $\sim 10.5$ seconds for 4000 posterior samples and so on. It should be remembered that this time is in addition to the time for cross-validation as well as the independent sampling from the conjugate posterior.
@@
The estimation errors for each of the models measured in terms of MSE are plotted in Figure 1 corresponding to the independent design case and in Figure 2 corresponding to the correlated design case. Although the MSE of the sparse projection-posterior method is higher in the independent design regime, it does not exceed the LASSO or the bootstrap LASSO. This is expected since the sparse projection posterior heavily relies on the choice of penalization. However, one should also note that in the correlated design case, the errors for the other models increased, thereby diminishing the gap in MSE among all competing methods. This showcases that our method performs robustly even when the covariates are correlated, which is actually more realistic. To see if this trend is maintained when correlation among variables is further increased, we conducted the simulation with the AR correlation structure for the predictors, but now with an enhanced correlation parameter $\rho=0.9$. Under such high correlation, the MSE of our method actually becomes lower than most of the competing methods. The boxplots capturing this phenomenon are given in Figure 3. Moreover, the proposed method performs robustly in terms of MSE when the true data generating process is non-normal (Table 5), whereas methods such as horseshoe, SSVS or SSLASSO incur very high MSEs.
Figure 1: Boxplots of the MSEs for different methods averaged over $M=100$ replications corresponding to the independent design case.
Figure 2: Boxplots of the MSEs for different methods averaged over $M=100$ replications corresponding to the correlated design case.
Figure 3: Boxplots of the MSEs for different methods averaged over $M=100$ replications corresponding to the highly correlated design case.
@@
Projection 1 0.084 (0.063) 0.964 (0.108) 1 0.018 (0.037) 0.991 (0.019)
$n=300$ $p=300$ $s_{0}=10$ LASSO 1 0.090 (0.016) 0.961 (0.104) 1 0.074 (0.088) 0.959 (0.048)
MCP 1 0.086 (0.015) 0.973 (0.093) 1 0.049 (0.085) 0.973 (0.047)
Horseshoe 1 0.010 (0.011) 0.986 (0.014) 1 0.008 (0.042) 0.984 (0.011)
SSVS 1 0 1 1 0 1
SSLASSO 1 0 1 1 0.001 (0.000) 0.999 (0.005)
Projection 1 0.023 (0.014) 0.978 (0.105) 1 0.051 (0.086) 0.972 (0.050)
$n=1000$ $p=100$ $s_{0}=10$ LASSO 1 0.091 (0.013) 0.965 (0.081) 1 0.129 (0.012) 0.979 (0.019)
MCP 1 0.052 (0.013) 0.974 (0.017) 1 0.108 (0.011) 0.996 (0.084)
Horseshoe 1 0.006 (0.023) 0.992 (0.013) 1 0.004 (0.034) 0.991 (0.007)
@@
In Table 2, we report the measures for variable selection corresponding to the small-$p$ ($n=1000,p=100$), same-$p$ ($n=300=p$) and large-$p$ ($n=100,p=2000$) regimes and both design types. The SSVS maintained its best selection performance in low- and high-dimensional cases, followed by the SSLASSO and the projection method. The model selection performance of the sparse projection posterior is robust under the correlated design setup. When $p<n$, the Bayesian LASSO selects many noise variables, causing its FDP to be the highest and, consequently, its MCC to be the lowest. The MCP does a good job of identifying the correct signals. On average, the sparse projection-posterior-based method is comparable to the horseshoe method. Overall, from this limited simulation study, it can be concluded that the proposed method has good selection properties. In addition, our method has good coverage and this makes our method preferable in comparison to the existing Bayesian methods. High-dimensional linear regression under sparsity is a very crowded domain and several ground-breaking research has been conducted in this area. Therefore, it is challenging to outperform these current approaches in all aspects; yet, our approach has an edge in terms of guaranteeing precise coverage.
@@
We compare the coverage of 95% confidence or credible regions from the different methods included in the simulation. In Tables 3 and 4, we report the mean coverage averaged over the $M$ replicates corresponding to the uncorrelated and correlated designs respectively, but separately for active and inactive variables. In general, coverage and length-wise, all methods do better when the variables are independent. On the Bayesian side, only the proposed sparse projection-posterior method reaches close to the desired coverage. The average lengths of all the intervals increase significantly to provide the same extent of coverage when the dimension $p$ grows with the sample size $n$. The existing Bayesian methods provide coverage if the parameter is zero but fail to attain 95% coverage for non-zero parameter values corresponding to active predictors. Also, the average coverage of signals deteriorates in the correlated design setting. However, when $p\gg n$, the only Bayesian method that provides good coverage of the signals is the sparse projection-posterior method. The plots in Figure 4 show the coverage and length of credible/confidence regions averaged over all the signal variables for all competing methods. For small-$p$ and $n=p$ cases, the sparse projection posterior method attains desirable coverage with a very small length, making it the best candidate. Its performance degrades when the dimension exceeds $n$ but is still better than most methods. We also show that the post-selection credible ellipsoid attains exact coverage under both independent and correlated designs. The scope to jointly quantify for the signal uncertainty helps avoid the debiasing map. Since debiasing is computationally heavy, bypassing it gives more computational edge. It should also be noted that the sparse projection posterior offers good to moderate coverage in very high dimensions even without debiasing.
@@
Debiased Projection 0.908 (0.094) 0.453 (0.023) 0.967 (0.004) 0.427 (0.003)
Projection 0.899 (0.013) 0.314 (0.034) 0.988 (0.006) 0.172 (0.002)
Credible Ellipsoid 0.853 (0.112) 0.344 (0.045) - -
@@
Horseshoe 0.941 (0.099) 0.102 (0.015) 0.999 (0.000) 0.205 (0.005)
Debiased Projection 0.945 (0.011) 0.159 (0.006) 0.977 (0.001) 0.267 (0.015)
Projection 0.940 (0.048) 0.301 (0.030) 0.997 (0.001) 0.129 (0.005)
Credible Ellipsoid 0.927 (0.008) 0.268 (0.025) - -
@@
Debiased Projection 0.966 (0.111) 0.103 (0.003) 0.949 (0.030) 0.088 (0.002)
Projection 0.942 (0.019) 0.333 (0.026) 0.998 (0.000) 0.126 (0.003)
Credible Ellipsoid 0.956 (0.016) 0.225 (0.010) - -
@@
Debiased Projection 0.868 (0.131) 0.695 (0.011) 0.982 (0.002) 0.548 (0.007)
Projection 0.847 (0.119) 0.311 (0.135) 0.992 (0.003) 0.151 (0.006)
Credible Ellipsoid 0.906 (0.052) 0.410 (0.002) - -
@@
Horseshoe 0.945 (0.107) 0.541 (0.029) 0.996 (0.001) 0.361 (0.007)
Debiased Projection 0.947 (0.013) 0.315 (0.010) 0.968 (0.007) 0.464 (0.007)
Projection 0.941 (0.031) 0.361 (0.056) 0.975 (0.002) 0.147 (0.004)
Credible Ellipsoid 0.941 (0.007) 0.338 (0.013) - -
@@
Debiased Projection 0.952 (0.101) 0.168 (0.002) 0.969 (0.019) 0.092 (0.008)
Projection 0.949 (0.042) 0.377 (0.013) 0.996 (0.001) 0.129 (0.002)
Credible Ellipsoid 0.940 (0.011) 0.211 (0.004) - -
@@
In Figure 5, we present individual coverage and interval lengths for two randomly selected signals and two randomly selected noise components in all four settings. These settings compare intervals produced by the horseshoe, sparse projection-posterior, debiased LASSO (ZnZ), and bootstrap LASSO for active (1 and 6) and noise (50 and 100) variables. Both signals have a true coefficient of 2. The plots show frequentist methods, and the Bayesian sparse posterior-projection method offers good coverage and comparable interval lengths. The horseshoe prior, in particular, exhibits slightly shorter lengths, especially for noise variables, consistent with results in Table 3. Importantly, our proposed method consistently covers the true signal. Finally in Table 5, we show how our method maintains good coverage under different non-normal error distributions.
@@
Performance
Metric
Error
Model
LASSO MCP
Bootstrap
LASSO
Horseshoe SSLASSO SSVS Projection
MSE Uniform 1.94 (0.262) 1.84 (0.151) 9.99 (2.85) 4.07 (2.15) 0.50 (1.89) 4.12 (2.88) 2.06 (0.376)
Laplace 5.43 (2.946) 2.49 (1.464) 5.17 (2.600) 8.20 (7.583) 2.94 (2.244) 4.27 (2.101) 5.66 (2.039)
Chi-square 4.36 (0.722) 18.71 (1.783) 6.26 (0.602) 16.62 (1.744) 12.58 (0.910) 5.10 (1.918) 3.52 (0.189)
MCC Uniform 0.926 (0.035) 0.996 (0.001) - 0.991 (0.001) 0.998 (0.000) 1 (0.000) 0.931 (0.017)
Laplace 0.924 (0.087) 0.977 (0.058) - 0.977 (0.037) 0.986 (0.023) 1 (0.000) 0.967 (0.0490
Chi-square 0.932 (0.066) 0.747 (0.113) - 0.874 (0.109) 0.606 (0.017) 0.706 (0.066) 0.934 (0.059)
Signal Coverage Uniform - - 0.88 (0.06) 0.91 (0.03) - 0.90 (0.01) 0.94 (0.02)
Laplace - - 0.86 (0.12) 0.90 (0.14) - 0.91 (0.07) 0.93 (0.13)
Chi-square - - 0.80 (0.05) 0.83 (0.15) - 0.70 (0.07) 0.78 (0.05)
Signal Length Uniform - - 0.692 (0.028) 0.771 (0.028) - 0.373 (0.021) 0.752 (0.011)
Laplace - - 0.832 (0.147) 0.982 (0.133) - 0.410 (0.013) 1.018 (0.010)
Chi-square - - 2.253 (0.164) 3.430 (0.295) - 2.880 (0.100) 1.781 (0.044)
Table 5: MSE, MCC, average signal coverage and average length of signal intervals reported for different methods with $M=100$ replications. MSEs are reported in the order of $10^{-4}$ with standard error reported in the order of $10^{-6}$.
4.2 Simulation Study 2: Distributed Computing